# Generalization of Diffusion Models Arises with a Balanced Representation Space

### v1

@@
Zekai Zhang, Xiao Li, Xiang Li, Lianghe Shi, Meng Wu, Molei Tao1, Qing Qu University of Michigan    $\cdot$    1Georgia Institute of Technology Joint first author Corresponding author Abstract Diffusion models excel at generating high-quality, diverse samples, yet they risk memorizing training data when overfit to the training objective. We analyze the distinctions between memorization and generalization in diffusion models through the lens of representation learning. By investigating a two-layer ReLU denoising autoencoder (DAE), we prove that: (i) memorization corresponds to the model storing raw training dataset in the learned weights for encoding and decoding, yielding localized, “spiky” representations; whereas (ii) generalization arises when the model captures local data statistics, producing “balanced” representations. Furthermore, we validate our theoretical findings on real-world unconditional and text-to-image diffusion models, demonstrating that the same representation structures emerge in deep generative models with significant practical implications. Building on these insights, we propose a representation-based method for detecting memorization and a training-free editing technique that allows precise control via representation steering. Together, our results highlight that learning good representations is central to novel and meaningful generative modelling. Keywords: Diffusion Models, Generalization, Representation Learning, Denoising AutoEncoders  Date: December 24, 2025
@@
Resources: To be released
@@ 1  Introduction
Diffusion models [ho2020denoising, lou2023discrete] have rapidly emerged as the dominant class of generative models, powering state-of-the-art systems such as Stable Diffusion [Rombach_2022_CVPR], Flux [labs2025flux1kontextflowmatching], and Veo [google2025veo3]. By iteratively denoising random noise, they achieve unprecedented scalability, controllability, and fidelity. However, their empirical success raises a fundamental question: in principle, the standard training objective (e.g., denoising score matching) admits a closed-form solution that merely memorizes training examples [yi2023generalization]; in practice, however, real-world models consistently produce novel and diverse outputs [zhang2024emergence, kadkhodaie2023generalization]. This distinct mismatch between theoretical expectation and observed behavior poses a critical gap in our understanding of diffusion model generalization, with direct implications for privacy, interpretability, and trustworthy deployment [somepalli2023diffusion].
@@
Addressing this question has drawn increasing attention in the machine learning community [zhang2024emergence, li2024understanding, wang2024diffusion, kadkhodaie2023generalization, gu2025memorization, bonnaire2025diffusion, zhang2025generalization, bertrand2025closed, zhang2025understanding], yet existing explanations remain far from satisfactory. Early works based on random feature models [li2023generalization, george2025denoising] provide useful insights but necessarily oversimplify model architectures. Analyses of linear models on Gaussian mixtures [li2024understanding, wang2024diffusion, wang2025analytical] shed light on generalization but cannot capture memorization. Another line of research explores inductive biases by constructing handcrafted closed-form solutions from empirical data to approximate U-Net performance [kamb2025an, niedoba2024towards, lukoianov2025locality, floros2025anisotropy], attributing success to principles such as locality and equivariance. While these advances are valuable, the findings remain fragmented and phenomenological, and a more unified account of how diffusion models both memorize and generalize is still lacking (see Section˜5 for a more detailed discussion of related work).
@@
To address these challenges, we develop a unified mathematical framework based on a theoretical analysis of a nonlinear two-layer ReLU denoising autoencoder (DAE) [vincent2011connection]. This framework not only unifies the characterization of memorization and generalization, but also bridges distribution learning with representation learning, offering profound practical implications. Specifically: (i) Memorization. We prove that when empirical samples are locally sparse, the network weights memorize and store individual training examples, leading to overfitting and hence memorization. (ii) Generalization. Conversely, when the empirical data are locally abundant, the weights effectively capture local data statistics, enabling the model to generate novel in-distribution samples.
@@
Crucially, our work provides a unique representation-centric perspective on generalization [tian2025provable], highlighting the pivotal role of bottleneck activations in DAE networks. This view is motivated by recent empirical evidence on the duality between distribution learning and representation learning in diffusion models [li2025understanding, xiang2025ddae++]: they inherently learn informative features for downstream tasks [kwon2022diffusion, chen2024deconstructing], and representation alignment regularization has been shown to accelerate training [yu2024representation]. Our theory makes this connection explicit: memorized samples are encoded as spiky activations concentrated on a few neurons, whereas generalized samples yield balanced representations that reflect the underlying distribution. These contrasting modes of representation learning manifest in distinct generation behaviors in terms of memorization or generalization, which we comprehensively validated across a range of models, including EDM [karras2022elucidating], Diffusion Transformers (DiT) [peebles2023scalable], and Stable Diffusion v1.4 [Rombach_2022_CVPR] (SD1.4).
@@ 2.1  A Denoising Perspective of Diffusion Models
Basics of diffusion models. Diffusion models comprise two processes: (i) a forward noising process and (ii) a reverse denoising/sampling process. The forward process progressively corrupts a clean sample $\bm{x}_{0}$ via $\bm{x}_{t}=\bm{x}_{0}+\sigma_{t}\bm{\epsilon}$ with $\bm{\epsilon}\sim\mathcal{N}(\bm{0},\bm{I})$, while the reverse process (e.g., DDIM [song2020denoising]) removes noise to generate data:
@@
where $\nabla\log p_{t}(\bm{x}_{t})$ is the score function of the marginal distribution of the noisy sample $\bm{x}_{t}$ at time $t$. To estimate $\nabla\log p_{t}(\bm{x}_{t})$, we use a denoising autoencoder (DAE) $\bm{f}_{\bm{\theta}}(\bm{x}_{t})$ [karras2022elucidating, li2025back] that predicts $\bm{x}_{0}$ from $\bm{x}_{t}$, so that
@@
where $\bm{f}_{\mathrm{gt}}(\bm{y}):=\mathbb{E}\!\left[\bm{x}\mid\bm{x}+\sigma_{t}\bm{\epsilon}=\bm{y};\,\bm{x}\sim p_{\mathrm{gt}}\right]$ is the ground-truth denoiser via Tweedie’s formula [efron2011tweedie]. Thus, the ideal (population) objective to learn the DAE is
@@
Generalization of diffusion models. In practice, we only have finitely many empirical samples $\bm{X}=\{\bm{x}_{i}\}_{i=1}^{n}$ with $\bm{x}_{i}\sim p_{\text{gt}}$. Accordingly, we work with the empirical distribution $p_{\mathrm{emp}}=\tfrac{1}{n}\sum_{i=1}^{n}\delta(\bm{x}-\bm{x}_{i})$, and Equation˜2 reduces to its empirical counterpart. Minimizing this empirical loss leads to the nonparametric empirical denoiser $\bm{f}_{\mathrm{emp}}$ [gu2025memorization], which maps a noisy input towards the nearest training samples:
@@
Sampling with $\bm{f}_{\mathrm{emp}}$ can provably reproduce training samples [zhang2024emergence, baptista2025memorization]. In practice, however, this empirical loss is minimized by taking the gradient descent over a parameterized neural network, which does not always overfit; instead, it can approximate the population denoiser $\bm{f}_{\mathrm{gt}}$ [niedoba2024towards]. In this paper, we aim to understand when a parameterized network overfits (learns $\bm{f}_{\mathrm{emp}}$) versus generalizes (learns $\bm{f}_{\mathrm{gt}}$).
@@ 2.2  Our Theoretical Framework
which is a standard approximation to data manifolds used in recent theoretical studies [wang2024diffusion, zhang2024emergence, li2025understanding, cui2023high, gatmiry2024learning, biroli2024dynamical, kamkari2024geometric, buchanan2025edgememorizationdiffusionmodels, li2025understanding, li2025towards].
Model parameterization and training loss. Following [vincent2011connection, chen2023score, zeno2023minimum, cui2025precise], we parameterize the DAE by a two-layer ReLU network:
@@
with $\bm{W}_{1},\bm{W}_{2}\in\mathbb{R}^{d\times p}$, $\cdot$$_{+}$ denoting ReLU and $\bm{h}(\cdot)$ stands for the representation. Training and sampling can be viewed as operating with a collection of DAEs across multiple noise levels. Following prior work [li2024understanding, zeno2025diffusion, zhang2024analyzing, hanfeature], we begin with a fixed noise level $\sigma$. The $\ell_{2}$-regularized training objective is
@@
We adopt (5) as a minimal, tractable model to analyze memorization and generalization. Recent works [lukoianov2025locality, li2024understanding, wang2024unreasonable] imply that real diffusion models exhibit approximate piecewise linearity; our ReLU model shares this structure and can be viewed as a local approximation of such networks. We verify this connection via an SVD analysis of denoiser Jacobians [kadkhodaie2023generalization, achilli2024losing] for EDM, SD1.4, and ReLU DAE in Appendix A.3: around generalized samples, the Jacobian reflects local data statistics as in Cor. 3.3, whereas around memorized samples it becomes noticeably low-rank and is dominated by the corresponding data vector, consistent with Cor. 3.2.
@@ 3  Main Theorems
Remarks. The proof is deferred to Section˜C.2. The local minimizer (7) consists of a block-wise main term plus a residual $\bm{R}(\sigma,\gamma)$, which vanishes as $\sigma$ becomes small relative to the separation margin $\gamma$. This is consistent with the low-noise regimes that are crucial for diffusion-model sampling and representation learning [niedoba2024towards, pavlova2025diffusion]. Empirically, we observe this block-wise structure even for relatively large $\sigma$ (Figure˜4). The $(\alpha,\beta)$-separability assumption serves mainly to simplify the proof; similar conclusions hold more generally (see Appendix A.1). Finally, the optimal solution is not tied to a specific block order, since $\bm{f}_{\bm{W}_{2},\bm{W}_{1}}$ is invariant to arbitrary column permutations of the weight matrices $(\bm{W}_{1},\bm{W}_{2})$.
@@ 3.1  Case 1: Memorization with Overparameterization
  • Learning the optimal solution with sparse columns. The structured solution with (pn)(p-n) trailing zero columns in (9) is one among many local minimizers, as dense alternatives can arise by splitting sparse columns. However, empirical evidence and theory [xie2024adam] suggest that standard optimizers such as Adam [kingma2014adam] bias training toward $\ell_{\infty}$-smooth solutions of the DAE loss (cf. Corollary˜C.4). As a result, the solutions observed in practice often align with the sparse structure we construct (Figure˜4).
  • Sampling reproduce training samples (memorization). In this regime, the learned DAE closely approximates the empirical denoiser $\bm{f}_{\mathrm{emp}}$ in Eq. (3), achieving low empirical loss and consequently reproducing the training samples under sampling (as shown in Figure˜3). This occurs because the DAE’s projection and reconstruction over the sparse columns of the weights during the reverse sampling effectively act as a power method, recovering memorized training data [weitzner2024relation]. Quantitatively, by plugging Corollary˜3.2 into the overall denoising score matching loss, we find that the KL divergence between the sampled and empirical distributions is bounded by $\frac{\pi}{2}\max_{1\leq i\leq n}\|\bm{x}_{i}\|$, confirming strong memorization.
@@
    This sparsity arises because $\bm{x}_{i}$ is negatively correlated with other samples stored in the learned weight matrix $\bm{W}_{\mathrm{mem}}$, yielding a nearly one-hot feature vector within the representation space (Figure˜5). Such spikiness could serve as a robust signature of memorization [hakemi2025deeper, gan2025massive], which we empirically demonstrate on both synthetic (Figure˜5) and real-world (Figure˜6) settings. Building on this insight, we introduce a simple yet effective memorization detection method that achieves strong results, as detailed later in Section˜4.1. Additionally, analogous correlations between sharp, localized activations and the recall of concrete stored knowledge have been empirically observed in Large Language Models (LLMs) [sun2024massive], suggesting our findings could also offer a potential explanation for these phenomena in LLMs.
@@ 3.2  Case 2: Generalization with Underparameterization
Under the problem setup of Theorem˜3.1, we assume the training data satisfy the separability condition in Definition˜3.1.111For simplicity, we take separability as an assumption; given sufficient samples, it can be verified under extra conditions on the means and covariances of MoG using standard measure concentration tools [Vershynin2018HDP]. If the DAE network in (5) is under-parameterized with $p=\sum_{k=1}^{K}p_{k}\ll n$, then there exists a local minimizer of the DAE training loss (6) such that
@@
  • Sampling yields novel in-distribution samples (generalization). When the model is underparameterized, our results show that the local optimal solution learned from the training data achieves bounded population loss on the MoG distribution by effectively acting as an optimal local denoiser for each mode. Consequently, sampling [li2024sharp, litowards] from the trained DAE produces in-distribution images that are distinct from the training samples, as illustrated in Figure˜3.
    Moreover, the population loss depends on the spectrum of $\bm{S}_{k}$ (equivalently, $\bm{\Sigma}_{k}$). When $\bm{\Sigma}_{k}$ has an approximately low-rank structure [de2022convergence, cole2024score, huang2024denoising], the loss is small and decays rapidly with the number of samples per mode $n_{k}$. This provides a principled explanation for the reproducibility of diffusion models across disjoint training subsets [zhang2024emergence, kadkhodaie2023generalization].
  • Balanced representations as a signature of generalization. Unlike the spiky representations in Corollary˜3.2, the underparameterized solution spreads the energy of $\bm{x}_{i}+\sigma\bm{\epsilon}$, with $\bm{x}_{i}\sim\mathcal{N}(\bm{\mu}_{k},\bm{\Sigma}_{k})$, across the $p_{k}$ coordinates of the active block (see Figure˜5). The representation behaves like a low-dimensional projection for a Gaussian mode [tipping1999probabilistic]:
@@
    Intuitively, generalized samples activate multiple neurons rather than a single spiky unit; the resulting projections encode information about the underlying distribution, helping to explain empirical findings on semantic directions [kwon2022diffusion] that are useful for editing, which we further explore in Section˜4.2.
Figure 6: Mem./Gen. representations in real-world models. memorized samples have spiky representations while generalized samples have more balanced ones. The layout follows Figure˜5 and the results are consistent with it. Representations are extracted at timestep $t=50$ ($\sigma_{t}\approx 0.17$). Left: DiT-L/4 pretrained on an ImageNet subset. Right: Stable Diffusion v1.4 pretrained on LAION [schuhmann2022laion]. Results for EDM pretrained on CIFAR10 and additional details are in Section˜B.2.
Concluding Corollary˜3.2 and Corollary˜3.3, we see the learned structure remains stable across timesteps, with $\sigma$ primarily acting as a regularization parameter. Varying $\sigma$ only slightly perturbs the solution, which helps explain the empirical success of diffusion models that employ a single neural network for denoising across multiple noise levels [sun2025noise].

### v2

@@
Zekai Zhang, Xiao Li, Xiang Li, Lianghe Shi, Meng Wu, Molei Tao1, Qing Qu University of Michigan    $\cdot$    1Georgia Institute of Technology Joint first author Corresponding author Abstract Diffusion models excel at generating high-quality, diverse samples, yet they risk memorizing training data when overfit to the training objective. We analyze the distinctions between memorization and generalization in diffusion models through the lens of representation learning. By investigating a two-layer ReLU denoising autoencoder (DAE), we prove that: (i) memorization corresponds to the model storing raw training dataset in the learned weights for encoding and decoding, yielding localized, spiky representations; whereas (ii) generalization arises when the model captures local data statistics, producing balanced representations. Furthermore, we validate our theoretical findings on real-world unconditional and text-to-image diffusion models, demonstrating that the same representation structures emerge in deep generative models with significant practical implications. Building on these insights, we propose a representation-based method for detecting memorization and a training-free editing technique that allows precise control via representation steering. Together, our results highlight that learning good representations is central to novel and meaningful generative modelling. Keywords: Diffusion Models, Generalization, Representation Learning, Denoising AutoEncoders  Date:
@@
Resources: Code Repository $\cdot$ Project Website
@@ 1  Introduction
Diffusion models [36, 56] have rapidly emerged as the dominant class of generative models, powering state-of-the-art systems such as Stable Diffusion [66], Flux [47], and Veo [30]. By iteratively denoising random noise, they achieve unprecedented scalability, controllability, and fidelity. However, their empirical success raises a fundamental question: in principle, the standard training objective (e.g., denoising score matching) admits a closed-form solution that merely memorizes training examples [103]; in practice, however, real-world models consistently produce novel and diverse outputs [110, 39]. This distinct mismatch between theoretical expectation and observed behavior poses a critical gap in our understanding of diffusion model generalization, with direct implications for privacy, interpretability, and trustworthy deployment [72].
@@
Addressing this question has drawn increasing attention in the machine learning community [110, 52, 89, 39, 31, 11, 109, 8, 108], yet existing explanations remain far from satisfactory. Early works based on random feature models [50, 29] provide useful insights but necessarily oversimplify model architectures. Analyses of linear models on Gaussian mixtures [52, 89, 87] shed light on generalization but cannot capture memorization. Another line of research explores inductive biases by constructing handcrafted closed-form solutions from empirical data to approximate U-Net performance [41, 60, 57, 25], attributing success to principles such as locality and equivariance. While these advances are valuable, the findings remain fragmented and phenomenological, and a more unified account of how diffusion models both memorize and generalize is still lacking (see Section˜5 for a more detailed discussion of related work).
@@
To address these challenges, we develop a unified mathematical framework based on a theoretical analysis of a nonlinear two-layer ReLU denoising autoencoder (DAE) [85]. This framework not only unifies the characterization of memorization and generalization, but also bridges distribution learning with representation learning, offering profound practical implications. Specifically: (i) Memorization. We prove that when empirical samples are locally sparse, the network weights memorize and store individual training examples, leading to overfitting and hence memorization. (ii) Generalization. Conversely, when the empirical data are locally abundant, the weights effectively capture local data statistics, enabling the model to generate novel in-distribution samples.
@@
Crucially, our work provides a unique representation-centric perspective on generalization [79], highlighting the pivotal role of bottleneck activations in DAE networks. This view is motivated by recent empirical evidence on the duality between distribution learning and representation learning in diffusion models [54, 97, 80]: they inherently learn informative features for downstream tasks [46, 18], and representation alignment regularization has been shown to accelerate training [104]. Our theory makes this connection explicit: memorized samples are encoded as spiky activations concentrated on a few neurons, whereas generalized samples yield balanced representations that reflect the underlying distribution. These contrasting modes of representation learning manifest in distinct generation behaviors in terms of memorization or generalization, which we comprehensively validated across a range of models, including EDM [43], Diffusion Transformers (DiT) [62], and Stable Diffusion v1.4 [66] (SD1.4).
@@ 2.1  A Denoising Perspective of Diffusion Models
Basics of diffusion models. Diffusion models comprise two processes: (i) a forward noising process and (ii) a reverse denoising/sampling process. The forward process progressively corrupts a clean sample $\bm{x}_{0}$ via $\bm{x}_{t}=\bm{x}_{0}+\sigma_{t}\bm{\epsilon}$ with $\bm{\epsilon}\sim\mathcal{N}(\bm{0},\bm{I})$, while the reverse process (e.g., DDIM [74]) removes noise to generate data:
@@
where $\nabla\log p_{t}(\bm{x}_{t})$ is the score function of the marginal distribution of the noisy sample $\bm{x}_{t}$ at time $t$. To estimate $\nabla\log p_{t}(\bm{x}_{t})$, we use a denoising autoencoder (DAE) $\bm{f}_{\bm{\theta}}(\bm{x}_{t})$ [43, 51] that predicts $\bm{x}_{0}$ from $\bm{x}_{t}$, so that
@@
where $\bm{f}_{\mathrm{gt}}(\bm{y}):=\mathbb{E}\!\left[\bm{x}\mid\bm{x}+\sigma_{t}\bm{\epsilon}=\bm{y};\,\bm{x}\sim p_{\mathrm{gt}}\right]$ is the ground-truth denoiser via Tweedie’s formula [24]. Thus, the ideal (population) objective to learn the DAE is
@@
Generalization of diffusion models. In practice, we only have finitely many empirical samples $\bm{X}=\{\bm{x}_{i}\}_{i=1}^{n}$ with $\bm{x}_{i}\sim p_{\text{gt}}$. Accordingly, we work with the empirical distribution $p_{\mathrm{emp}}=\tfrac{1}{n}\sum_{i=1}^{n}\delta(\bm{x}-\bm{x}_{i})$, and Equation˜2 reduces to its empirical counterpart. Minimizing this empirical loss leads to the nonparametric empirical denoiser $\bm{f}_{\mathrm{emp}}$ [31], which maps a noisy input towards the nearest training samples:
@@
Sampling with $\bm{f}_{\mathrm{emp}}$ can provably reproduce training samples [110, 5]. In practice, however, this empirical loss is minimized by taking the gradient descent over a parameterized neural network, which does not always overfit; instead, it can approximate the population denoiser $\bm{f}_{\mathrm{gt}}$ [60]. In this paper, we aim to understand when a parameterized network overfits (learns $\bm{f}_{\mathrm{emp}}$) versus generalizes (learns $\bm{f}_{\mathrm{gt}}$).
@@ 2.2  Our Theoretical Framework
which is a standard approximation to data manifolds used in recent theoretical studies [89, 110, 54, 22, 28, 10, 42, 12, 54, 53].
Model parameterization and training loss. Following [85, 16, 106, 21], we parameterize the DAE by a two-layer ReLU network:
@@
with $\bm{W}_{1},\bm{W}_{2}\in\mathbb{R}^{d\times p}$, $\cdot$$_{+}$ denoting ReLU and $\bm{h}(\cdot)$ stands for the representation. Training and sampling can be viewed as operating with a collection of DAEs across multiple noise levels. Following prior work [52, 105, 107, 33], we begin with a fixed noise level $\sigma$. The $\ell_{2}$-regularized training objective is
@@
We adopt (5) as a minimal, tractable model to analyze memorization and generalization. Recent works [57, 52, 86] imply that real diffusion models exhibit approximate piecewise linearity; our ReLU model shares this structure and can be viewed as a local approximation of such networks. We verify this connection via an SVD analysis of denoiser Jacobians [39, 1] for EDM, SD1.4, and ReLU DAE in Appendix A.3: around generalized samples, the Jacobian reflects local data statistics as in Cor. 3.3, whereas around memorized samples it becomes noticeably low-rank and is dominated by the corresponding data vector, consistent with Cor. 3.2.
@@ 3  Main Theorems
Remarks. The proof is deferred to Section˜C.2. The local minimizer (7) consists of a block-wise main term plus a residual $\bm{R}(\sigma,\gamma)$, which vanishes as $\sigma$ becomes small relative to the separation margin $\gamma$. This is consistent with the low-noise regimes that are crucial for diffusion-model sampling and representation learning [60, 61]. Empirically, we observe this block-wise structure even for relatively large $\sigma$ (Figure˜4). The $(\alpha,\beta)$-separability assumption serves mainly to simplify the proof; similar conclusions hold more generally (see Appendix A.1). Finally, the optimal solution is not tied to a specific block order, since $\bm{f}_{\bm{W}_{2},\bm{W}_{1}}$ is invariant to arbitrary column permutations of the weight matrices $(\bm{W}_{1},\bm{W}_{2})$.
@@ 3.1  Case 1: Memorization with Overparameterization
  • Learning the optimal solution with sparse columns. The structured solution with (pn)(p-n) trailing zero columns in (9) is one among many local minimizers, as dense alternatives can arise by splitting sparse columns. However, empirical evidence and theory [99] suggest that standard optimizers such as Adam [44] bias training toward $\ell_{\infty}$-smooth solutions of the DAE loss (cf. Corollary˜C.4). As a result, the solutions observed in practice often align with the sparse structure we construct (Figure˜4).
  • Sampling reproduce training samples (memorization). In this regime, the learned DAE closely approximates the empirical denoiser $\bm{f}_{\mathrm{emp}}$ in Eq. (3), achieving low empirical loss and consequently reproducing the training samples under sampling (as shown in Figure˜3). This occurs because the DAE’s projection and reconstruction over the sparse columns of the weights during the reverse sampling effectively act as a power method, recovering memorized training data [93]. Quantitatively, by plugging Corollary˜3.2 into the overall denoising score matching loss, we find that the KL divergence between the sampled and empirical distributions is bounded by $\frac{\pi}{2}\max_{1\leq i\leq n}\|\bm{x}_{i}\|$, confirming strong memorization.
@@
    This sparsity arises because $\bm{x}_{i}$ is negatively correlated with other samples stored in the learned weight matrix $\bm{W}_{\mathrm{mem}}$, yielding a nearly one-hot feature vector within the representation space (Figure˜5). Such spikiness could serve as a robust signature of memorization [32, 26], which we empirically demonstrate on both synthetic (Figure˜5) and real-world (Figure˜6) settings. Building on this insight, we introduce a simple yet effective memorization detection method that achieves strong results, as detailed later in Section˜4.1. Additionally, analogous correlations between sharp, localized activations and the recall of concrete stored knowledge have been empirically observed in Large Language Models (LLMs) [77], suggesting our findings could also offer a potential explanation for these phenomena in LLMs.
@@ 3.2  Case 2: Generalization with Underparameterization
Under the problem setup of Theorem˜3.1, we assume the training data satisfy the separability condition in Definition˜3.1.111For simplicity, we take separability as an assumption; given sufficient samples, it can be verified under extra conditions on the means and covariances of MoG using standard measure concentration tools [83]. If the DAE network in (5) is under-parameterized with $p=\sum_{k=1}^{K}p_{k}\ll n$, then there exists a local minimizer of the DAE training loss (6) such that
@@
  • Sampling yields novel in-distribution samples (generalization). When the model is underparameterized, our results show that the local optimal solution learned from the training data achieves bounded population loss on the MoG distribution by effectively acting as an optimal local denoiser for each mode. Consequently, sampling [49, 48] from the trained DAE produces in-distribution images that are distinct from the training samples, as illustrated in Figure˜3.
    Moreover, the population loss depends on the spectrum of $\bm{S}_{k}$ (equivalently, $\bm{\Sigma}_{k}$). When $\bm{\Sigma}_{k}$ has an approximately low-rank structure [23, 20, 37], the loss is small and decays rapidly with the number of samples per mode $n_{k}$. This provides a principled explanation for the reproducibility of diffusion models across disjoint training subsets [110, 39].
  • Balanced representations as a signature of generalization. Unlike the spiky representations in Corollary˜3.2, the underparameterized solution spreads the energy of $\bm{x}_{i}+\sigma\bm{\epsilon}$, with $\bm{x}_{i}\sim\mathcal{N}(\bm{\mu}_{k},\bm{\Sigma}_{k})$, across the $p_{k}$ coordinates of the active block (see Figure˜5). The representation behaves like a low-dimensional projection for a Gaussian mode [81]:
@@
    Intuitively, generalized samples activate multiple neurons rather than a single spiky unit; the resulting projections encode information about the underlying distribution, helping to explain empirical findings on semantic directions [46] that are useful for editing, which we further explore in Section˜4.2.
Figure 6: Mem./Gen. representations in real-world models. memorized samples have spiky representations while generalized samples have more balanced ones. The layout follows Figure˜5 and the results are consistent with it. Representations are extracted at timestep $t=50$ ($\sigma_{t}\approx 0.17$). Left: DiT-L/4 pretrained on an ImageNet subset. Right: Stable Diffusion v1.4 pretrained on LAION [69]. Results for EDM pretrained on CIFAR10 and additional details are in Section˜B.2.
Concluding Corollary˜3.2 and Corollary˜3.3, we see the learned structure remains stable across timesteps, with $\sigma$ primarily acting as a regularization parameter. Varying $\sigma$ only slightly perturbs the solution, which helps explain the empirical success of diffusion models that employ a single neural network for denoising across multiple noise levels [78].